Question 1) (60 Marks) a) Given the joint pdf of the continuous RVs X and Y: fxy(x, y) = c for the region {0 sxs y, 0 sy s 1}  

need urgent, no need to write it neat and clean, just send solutions in 10 to 15 minutes..

b0bGr..
Answer #1

a) The point pdf is O느거느y의 finy) = e . We know that SS fan de daal. xy SO, is e dy dx = 1 gl dn=1 or, selym dn = 1 on s c(1-x6 Let de fine events and s represent illness and Smokep Here given values are P(11) - 0.1 P(IIS) = 0.005 -P(S)=20% = 0.20 P (